OpenStudy (anonymous):

About 55% of emergency room visits are unnecessary. Assuming each emergency room visit is independent. (Hints: first define the random variable X. Then write the objective of each question in terms of X.) a) Suppose the distribution of unnecessary emergency room visits can be approximated by a normal distribution. What is the probability that AT LEAST two of them need to be in the emergency room? @dumbcow

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Binomial_distribution#Normal_approximation what is sample size?

OpenStudy (anonymous):

15 patients

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

mean = .55 std dev = sqrt(15*.55*.45) \[P(X>2/15) = 1- P(X< 2/15)\] \[Z = \frac{2/15 - 0.55}{\sqrt{15*.55*.45}}\] look up probability using z-table

OpenStudy (anonymous):

is it 2 divided by 15?

OpenStudy (anonymous):

@dumbcow

OpenStudy (amistre64):

P(0)+P(1) is prolly an equivalent way to approach it

OpenStudy (amistre64):

0 needs + 1 needs; .55 not, .45 needs \[\binom{15}{0}.45^0.55^{15}+\binom{15}{1}.45^1.55^{14}\]

OpenStudy (amistre64):

1 - that

OpenStudy (anonymous):

oh ok...this much understandable..do i subtract my final answer from 1?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

oh nice

OpenStudy (anonymous):

thank you..very much..

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

@amistre64 WHY doesnt .55 come before .45 in both P(0) + P(1)

OpenStudy (amistre64):

About 55% of emergency room visits are unnecessary. .55 = not need What is the probability that AT LEAST two of them NEED to be in the emergency room? exactly 0 need, exactly 1 need, ex 2 need, ex 3 need, ..., 15 need |-------------------------| NOT "at least 2 need" 1 - (not at least 2 need) = (at least 2 need) since we are looking for the prolly of NEED as a success, 1-not need = need 1 - .55 = .45 the setup is: (n k) s^k f^(n-k) P(0 need) = (15 0) .45^0 .55^15 P(1 need) = (15 1) .45^1 .55^14